# No-Regret Strategy Solving in Imperfect-Information Games via Pre-Trained Embedding

### v1

@@ Abstract
High-quality information set abstraction remains a core challenge in solving large-scale imperfect-information extensive-form games (IIEFGs) — such as no-limit Texas Hold’em — where the finite nature of spatial resources hinders strategy solving over the full game. State-of-the-art AI methods rely on pre-trained discrete clustering for abstraction, yet their hard classification irreversibly loses critical information: specifically, the quantifiable subtle differences between information sets — vital for strategy solving — thereby compromising the quality of such solving. Inspired by the word embedding paradigm in natural language processing, this paper proposes the Embedding CFR algorithm, a novel approach for solving strategies in IIEFGs within an embedding space. The algorithm pre-trains and embeds features of isolated information sets into an interconnected low-dimensional continuous space, where the resulting vectors more precisely capture both the distinctions and connections between information sets. Embedding CFR presents a strategy-solving process driven by regret accumulation and strategy updates within this embedding space, with accompanying theoretical analysis verifying its capacity to reduce cumulative regret. Experiments on poker show that with the same spatial overhead, Embedding CFR achieves significantly faster exploitability convergence compared to cluster-based abstraction algorithms, confirming its effectiveness. Furthermore, to our knowledge, it is the first algorithm in poker AI that pre-trains information set abstractions through low-dimensional embedding for strategy solving.
@@ 1 Introduction
Solving large-scale imperfect-information games like no-limit Texas Hold’em remains a pivotal challenge in AI research. Systems including DeepStack (Moravčík et al. 2017), Libratus (Brown and Sandholm 2018), and Pluribus (Brown and Sandholm 2019b) have achieved superhuman performance through game-theoretic equilibrium computation, yet their success hinges on addressing a critical dilemma: handling enormous decision spaces with limited spatial resources.
Counterfactual Regret Minimization (CFR) serves as the foundational algorithm for computing $\epsilon$-Nash equilibrium, but its linear space complexity becomes prohibitive even for moderate-scale games. Consider head-up limit Texas Hold’em: with $\approx 10^{13}$ information sets, it demands 523 terabytes of RAM (Johanson 2013)—a scale rendering direct CFR implementation physically impossible, let alone larger games like head-up no-limit. Existing systems thus trade theoretical rigor for practicality via approximations, with the prevailing “abstraction-solving-translation” paradigm (Gilpin, Sandholm, and Sørensen 2007) as the core approach: compressing the original game into a feasible abstract version, solving for $\epsilon$-Nash equilibrium within it, and mapping the solved strategy back to the original game. One key approach within this paradigm, known as information set abstraction, is a pre-training method that groups similar information sets into discrete equivalence classes prior to strategy solving where uniform strategies are applied to these classes to reduce spatial resource demands. This approach is validated by the aforementioned superhuman AI systems, which leveraged it to defeat top human players.
However, existing information set abstraction algorithms have critical limitations. Relying predominantly on many-to-one mappings, they often introduce arbitrary classification boundaries. As shown in Figure 1(c), consider two hands in the flop round of Texas Hold’em (see Appendix A.1 for rules): hand $\blacksquare$ ($\texttt{9}^{p}_{s}\texttt{A}^{p}_{s}\texttt{9}_{h}\texttt{J}_{s}\texttt{Q}_{s}$)111h: hearts, d: diamonds, s: spades, c: clubs; superscripts p denote private cards (e.g., $\texttt{8}^{p}_{h}\texttt{8}^{p}_{d}$= player’s 8s in hearts/diamonds). and hand $\bullet$ ($\texttt{T}^{p}_{s}\texttt{A}^{p}_{s}\texttt{T}_{h}\texttt{J}_{s}\texttt{Q}_{s}$); both hands forming a pair. While superficially similar, hand $\bullet$ can form a straight flush with $\texttt{K}_{s}$, making it inherently stronger than hand $\blacksquare$ (which can realistically only form a flush with any additional spade card). Current algorithms face a dilemma: keeping these hands — and the information sets they represent — separate wastes the similar strategies that would naturally follow from their inherent similarities, while grouping them overlooks nuanced differences. This binary, low-granularity approach introduces arbitrariness that degrades strategy quality.
@@
(c) Traditional abstraction maps information sets to a fixed number of $m$ abstract classes, e.g. buckets, forcing binary decisions for these hands: either refining $\blacksquare$ and
@@
In this paper, we propose Embedding CFR, a novel strategy-solving paradigm for information set abstraction, to address the aforementioned limitations. Unlike traditional methods that map multiple information sets to a single equivalence class, Embedding CFR maps each information set to a multi-dimensional vector—embedding coordinates that form a probability distribution — thereby enhancing the expressive power of information set abstraction. As illustrated in Figure 1(a), hands $\blacksquare$, $\bullet$ , and $\blacksquare$ are assigned embedding coordinates, where hand $\blacksquare$ ($\texttt{K}^{p}_{s}\texttt{A}^{p}_{s}\texttt{T}_{s}\texttt{J}_{s}\texttt{Q}_{s}$) forms a powerful straight flush at the flop. The coordinates of these hands in the embedding space represent their confidence in corresponding dimensions. During strategy solving, updates to information sets preferentially influence high-confidence dimensions, ensuring strategies associated with closer coordinates exhibit greater similarity. Figure 1(b) shows a schematic 2D projection of this embedding, where $\blacksquare$ and $\bullet$ cluster closely yet remain subtly distinguishable — yielding correspondingly similar solved strategies — while being distinctly separated from the strong hand $\blacksquare$ , ensuring minimal mutual influence during solving.
@@
Experiments are conducted in a poker game, comparing with traditional information set abstraction algorithms such as EHS (Gilpin and Sandholm 2007a), PaEmd (the information set abstraction algorithm adopted by Libratus, recognized as a state-of-the-art approach in the field) (Ganzfried and Sandholm 2014), and a recent work KrwEmd (Fu et al. 2024). The results show that the strategy solved by Embedding CFR has lower exploitability under the same spatial overhead and update iterations, demonstrating the effectiveness of the proposed algorithm.
@@ 1.1 Related Works
Our work lies within the community of solving imperfect-information extensive-form games via no-regret learning, with CFR (Zinkevich et al. 2007) as the core strategy-solving algorithm. Extensions follow two distinct directions: sampling-based methods like MCCFR (Lanctot et al. 2009), and strategy update-focused optimizations including CFR+ (Bowling et al. 2015), DCFR (Brown and Sandholm 2019a), and PCFR+ (Farina, Kroer, and Sandholm 2021).
For strategy solving under limited space constraints, related efforts include pruning and action abstraction techniques: regret-based pruning (Brown and Sandholm 2015), best-response pruning (Brown and Sandholm 2017), dynamic thresholding pruning (Brown, Kroer, and Sandholm 2017), along with reinforcement learning-based online action abstraction methods like RL-CFR (Li, Fang, and Huang 2024) and EVPA (Li and Huang 2025).
Closer to our approach, neural network-based regret estimation methods (DeepStack (Moravčík et al. 2017), ReBeL (Brown et al. 2020), Deep CFR (Brown et al. 2019), Escher (McAleer et al. 2023)) replace tabular storage of regrets and strategies to simplify solving. They excel in model-free scenarios, where reliance on environmental interactions — which simultaneously drive strategy optimization and information set learning — reduces dependence on complex domain knowledge. In model-known settings, however, this dual reliance creates trade-offs, blunting efficiency compared to pre-training-based methods leveraging prior domain knowledge.
Our closest predecessors are information set abstraction works: lossless abstraction applicable to small-scale games (Gilpin and Sandholm 2007b), expectation-based EHS (Gilpin and Sandholm 2007a), potential-aware methods (Gilpin, Sandholm, and Sørensen 2007; Ganzfried and Sandholm 2014) (notably PaEmd), and the history trajectory-based KrwEmd (Fu et al. 2024). We share with these predecessors a pre-training-then-solving paradigm — analogous to NLP pipelines where word embedding or tokenization precedes semantic processing — by first preprocessing information sets before strategy solving.
@@ 2 Background and Notation
For a given infoset $I$, we define $Z_{I}$ as the set of terminal histories that can be reached from some $h\in I$. Throughout this work, we assume the game adheres to the property of perfect recall, which stipulates that players retain complete knowledge of all information they have previously encountered. More formally, for player $i$, if two different histories are not part of the same infoset, then no subsequent continuations of these histories can be grouped within the same infoset for that player. An important implication of this property is that each branch from the root to the leaves of the game tree traverses any given infoset at most once. We use the notation $z[I]$ to denote the history prefix of a terminal history $z$ that passes through infoset $I$. This notation is undefined (or invalid) if the path from $h^{0}$ to $z$ does not traverse $I$.
@@
$\pi_{i}^{\sigma}(h,h^{\prime})=\begin{cases}0,&\text{if }h\not\sqsubseteq h^{\prime}\\ \prod\limits_{h\sqsubseteq h^{\prime\prime},h^{\prime\prime}\cdot a\sqsubseteq h^{\prime},\mathcal{P}(h^{\prime\prime})=i}\sigma_{i}(h^{\prime\prime},a),&\text{if }h\sqsubseteq h^{\prime}\end{cases}.$
@@
In game theory, the Nash equilibrium is a crucial solution concept. It represents a state where no player can enhance their payoff by unilaterally altering their strategy, and to a certain extent, it can be regarded as the optimal solution. The payoff for player $i$ under strategy profile $\sigma$ is $u_{i}(\sigma)=\sum_{z\in Z}u_{i}(z)\pi^{\sigma}(z)$. The best response value for player $i$ against $\sigma_{-i}$ is $b_{i}(\sigma_{-i})=\max_{\sigma_{i}^{\prime}\in\Sigma_{i}}u_{i}(\sigma_{i}^{\prime},\sigma_{-i})$, the max payoff for optimal response. A strategy profile $\sigma^{*}$ is an $\epsilon$-Nash equilibrium if $\forall i\in\mathcal{N}\backslash\{c\}$, $u_{i}(\sigma^{*})+\epsilon\geq b_{i}(\sigma^{*}_{-i})$. When $\epsilon=0$, it’s a standard Nash equilibrium. In two-player zero-sum games, the exploitability of $\sigma$ is $\epsilon^{\sigma}=b_{1}(\sigma_{2})+b_{2}(\sigma_{1})$, measuring players’ vulnerability to an optimal opponent. Lower exploitability means closer to equilibrium.
@@ 2.1 Counterfactual Regret Minimization
Counterfactual Regret Minimization (CFR), also referred to as vanilla CFR, is an iterative algorithm for computing approximate Nash equilibrium in IIEFGs by minimizing counterfactual regret (Zinkevich et al. 2007). For player $i$, the total regret after $T$ iterations is defined as
@@
In two-player zero-sum IIEFGs, if both players satisfy $R^{T}_{i}\leq\epsilon$, then the average strategy forms a $2\epsilon$-Nash equilibrium (Waugh 2009).
@@
is used to update the immediate strategy via regret matching (Hart and Mas-Colell 2000):
$\sigma^{T+1}_{\mathcal{P}(I)}(I,a)=\begin{cases}\frac{R^{T}(I,a)_{+}}{~~\sum\limits_{\mathclap{a^{\prime}\in A(I_{q})}}\left(R^{T}(I,a^{\prime})_{+}\right)},&\text{if }\sum\limits_{\mathclap{a^{\prime}\in A(I_{q})}}\left(R^{T}(I,a^{\prime})_{+}\right)>0,\\ \frac{1}{|A(I)|},&\text{otherwise,}\end{cases}$
@@ 3.2 Driving Process for Embedding CFR
Building on this notation, we propose Embedding CFR, which models the extended information set abstraction problem as a strategy voting framework, where a embedding matrix $\mathbf{\Phi}=[\phi]_{m\times n}$ is constructed. Each non-negative element $\phi_{p,q}\geq 0$ — with the constraint that $\sum_{p=1}^{m}\phi_{p,q}=1$ for each $q$ — represents the confidence in trusting advisor $e_{p}$ at infoset $I_{q}$; here, $\mathbf{\Phi}_{\cdot,q}$ denotes the embedding coordinates of infoset $I_{q}$. Given a strategy $\sigma_{i}^{T}(\mathbf{E})$ in the advisor space (also called embedding space), decisions in the original strategy space are made via the mapping
@@
To solve games under this abstraction, Embedding CFR operates as a no-regret learning algorithm that draws insights from vanilla CFR (Zinkevich et al. 2007): as illustrated in the right half of Figure 2 and as we have detailed in Section 2.1, vanilla CFR updates the cumulative counterfactual regret $R^{T}(I)$ of the target infoset $I$ at iteration $T$, then derives the immediate strategy $\sigma_{i}^{T+1}(I)$ for next-iteration online interaction with the environment via regret matching, after which the regret for the new iteration is updated—forming a driving process. The strategy to be evaluated, which exhibits favorable convergence properties, is the average strategy $\bar{\sigma}_{i}^{T}(I)$. Embedding CFR extends this process by addressing how to drive it within the newly introduced advisor space.
@@
It is important to note that while regret embedding (Equation (5)) and strategy recovery (Equations (4), (8)) involve full matrix multiplications — specifically of sizes $m\times n$ with $n\times m$ and $n\times m$ with $m\times n$ — this might initially suggest unavoidable $\mathcal{O}(n)$ or higher space complexity. Our solution addresses this by framing regret embedding as a sampling process and strategy recovery as a query procedure. Rather than operating on the full set of infosets, we dynamically engage a subset of $l<n$ infosets $I_{(1)},\dots,I_{(l)}$ (mapping to indices $j_{1},\dots,j_{l}$ in $\mathbf{J}$) at each iteration. As visualized in Figure 2, this selective focus is concretely illustrated: iteration $T$ activates only the single infoset $I_{4}$, while iteration $T+1$ shifts to $I_{3}$.
@@
An additional consideration is how to efficiently retrieve the embedding coordinates $\mathbf{\Phi}_{\cdot,l_{k}}$ for each infoset $I_{(k)}$ — this will be discussed in the next section. Notably, each $\mathbf{\Phi}_{\cdot,l_{k}}$ only incurs $\mathcal{O}(m)$ space overhead, which does not significantly add to overall storage requirements. The blue and green arrows in Figure 2 illustrate these dynamic pathways, with purple nodes representing the quantities that need to be stored in this iteration—emphasizing that this paired sampling-and-query mechanism ensures Embedding CFR operates with $\mathcal{O}(m)$ space complexity.

### v2

@@ Abstract
High-quality information set abstraction remains a core challenge in solving large-scale imperfect-information extensive-form games (IIEFGs)–such as no-limit Texas Hold’em–where the finite nature of spatial resources hinders solving strategies for the full game. State-of-the-art AI methods rely on pre-trained discrete clustering for abstraction, yet their hard classification irreversibly discards critical information: specifically, the quantifiable subtle differences between information sets—vital for strategy solving—thus compromising the quality of such solving. Inspired by the word embedding paradigm in natural language processing, this paper proposes the Embedding CFR algorithm, a novel approach for solving strategies in IIEFGs within an embedding space. The algorithm pre-trains and embeds the features of individual information sets into an interconnected low-dimensional continuous space, where the resulting vectors more precisely capture both the distinctions and connections between information sets. Embedding CFR introduces a strategy-solving process driven by regret accumulation and strategy updates in this embedding space, with supporting theoretical analysis verifying its ability to reduce cumulative regret. Experiments on poker show that with the same spatial overhead, Embedding CFR achieves significantly faster exploitability convergence compared to cluster-based abstraction algorithms, confirming its effectiveness. Furthermore, to our knowledge, it is the first algorithm in poker AI that pre-trains information set abstractions via low-dimensional embedding for strategy solving.
Codehttps://github.com/PhilEnchan/EmbeddingCFR
@@ 1 Introduction
Solving large-scale imperfect-information games like no-limit Texas Hold’em remains a pivotal challenge in AI research. Systems including DeepStack (moravvcik2017deepstack), Libratus (brown2018superhuman), and Pluribus (brown2019superhuman) have achieved superhuman performance through game-theoretic equilibrium computation, yet their success hinges on addressing a critical dilemma: handling enormous decision spaces with limited spatial resources.
Counterfactual Regret Minimization (CFR) serves as the foundational algorithm for computing $\epsilon$-Nash equilibrium, but its linear space complexity becomes prohibitive even for moderate-scale games. Consider heads-up limit Texas Hold’em: with $\approx 10^{13}$ information sets, it demands 523 terabytes of RAM (johanson2013measuring)—a scale rendering direct CFR implementation physically impossible, let alone larger games like heads-up no-limit. Existing systems thus trade theoretical rigor for practicality via approximations, with the prevailing “abstraction-solving-translation” paradigm (gilpin2007potential) as the core approach: compressing the original game into a feasible abstract version, solving for $\epsilon$-Nash equilibrium within it, and mapping the solved strategy back to the original game. One key approach within this paradigm, known as information set abstraction, is a pre-training method that groups similar information sets into discrete equivalence classes prior to strategy solving, and uniform strategies are applied to these classes to reduce spatial resource demands. This approach is validated by the aforementioned superhuman AI systems, which leveraged it to defeat top human players.
However, existing information set abstraction algorithms have critical limitations. Relying predominantly on many-to-one mappings, they often introduce arbitrary classification boundaries. As shown in Figure 1(c), consider two hands in the flop round of Texas Hold’em (see Appendix A.1 for rules) [^1] All appendix content is available in the extended version.: hand $\blacksquare$ ($\texttt{9}^{p}_{s}\texttt{A}^{p}_{s}\texttt{9}_{h}\texttt{J}_{s}\texttt{Q}_{s}$)222h: hearts, d: diamonds, s: spades, c: clubs; superscripts p denote private cards (e.g., $\texttt{8}^{p}_{h}\texttt{8}^{p}_{d}$= player’s 8s in hearts/diamonds). and hand $\bullet$ ($\texttt{T}^{p}_{s}\texttt{A}^{p}_{s}\texttt{T}_{h}\texttt{J}_{s}\texttt{Q}_{s}$); both hands forming a pair. While superficially similar, hand $\bullet$ can form a straight flush with $\texttt{K}_{s}$, making it inherently stronger than hand $\blacksquare$ (which can realistically only form a flush with any additional spade card). Current algorithms face a dilemma: keeping these hands–and the information sets they represent–separate wastes the similar strategies that would naturally follow from their inherent similarities, while grouping them overlooks nuanced differences. This binary, low-granularity approach introduces arbitrariness that degrades strategy quality.
@@
(c) Traditional abstraction maps information sets to a fixed number of $m$ abstracted classes, e.g. buckets, forcing binary decisions for these hands: either refining $\blacksquare$ and
@@
In this paper, we propose Embedding CFR, a novel strategy-solving paradigm for information set abstraction, to address the aforementioned limitations. Unlike traditional methods that map multiple information sets to a single equivalence class, it maps each information set to a multi-dimensional vector—embedding coordinates forming a probability distribution–thereby enhancing abstraction expressiveness. As Figure 1(a) illustrates, hands $\blacksquare$, $\bullet$ , and $\blacksquare$ (the latter a strong straight flush: $\texttt{K}^{p}_{s}\texttt{A}^{p}_{s}\texttt{T}_{s}\texttt{J}_{s}\texttt{Q}_{s}$ at the flop) are assigned embedding coordinates representing their confidence in corresponding dimensions. During strategy solving, updates to information sets preferentially influence high-confidence dimensions, ensuring strategies associated with closer coordinates exhibit greater similarity. Figure 1(b) shows a schematic 2D projection of this embedding, where hand $\blacksquare$ and hand $\bullet$ cluster closely yet remain subtly distinguishable–yielding correspondingly similar solved strategies–while being distinctly separated from the strong hand $\blacksquare$ , ensuring minimal mutual influence during solving.
@@
Experiments are conducted in a poker game, comparing with traditional information set abstraction algorithms such as EHS (gilpin2007better), PaEmd (the information set abstraction algorithm adopted by Libratus, recognized as a state-of-the-art approach in the field) (ganzfried2014potential), and a recent work KrwEmd (fu2024signal). The results show that the strategy solved by Embedding CFR has lower exploitability under the same spatial overhead and update iterations, demonstrating the effectiveness of the proposed algorithm.
@@ 1.1 Related Works
Our work lies within the community of solving imperfect-information extensive-form games via no-regret learning, with CFR (zinkevich2007regret) as the core strategy-solving algorithm. Extensions follow two distinct directions: sampling-based methods like MCCFR (lanctot2009monte), and strategy update-focused optimizations including CFR+ (bowling2015heads), DCFR (brown2019solving), and PCFR+ (farina2021faster).
For strategy solving under limited space constraints, related efforts include pruning and action abstraction techniques: regret-based pruning (brown2015regret), best-response pruning (brown2017reduced), dynamic thresholding pruning (brown2017dynamic), along with reinforcement learning-based online action abstraction methods like RL-CFR (li2024rl) and EVPA (li2025efficient).
Closer to our approach, neural network-based regret estimation methods (DeepStack (moravvcik2017deepstack), ReBeL (brown2020combining), Deep CFR (brown2019deep), Escher (mcaleer2023escher)) replace tabular storage of regrets and strategies to simplify solving. They excel in model-free scenarios, where reliance on environmental interactions–which simultaneously drive strategy optimization and information set learning–reduces dependence on complex domain knowledge. In model-known settings, however, this dual reliance creates trade-offs, blunting efficiency compared to pre-training-based methods leveraging prior domain knowledge.
Our closest predecessors are information set abstraction works: lossless abstraction applicable to small-scale games (gilpin2007lossless), expectation-based EHS (gilpin2007better), potential-aware methods (gilpin2007potential; ganzfried2014potential) (notably PaEmd), and the history trajectory-based KrwEmd (fu2024signal). We share with these predecessors a pre-training-then-solving paradigm–analogous to NLP pipelines where word embedding or tokenization precedes semantic processing–by first preprocessing information sets before strategy solving.
@@ 2 Background and Notation
For infoset $I$, we define $Z_{I}$ as the set of terminal histories that can be reached from some $h\in I$. Throughout this work, we assume the game adheres to the property of perfect recall, which stipulates that players retain complete knowledge of all information they have previously encountered. More formally, for player $i$, if two different histories are not part of the same infoset, then no subsequent continuations of these histories can be grouped within the same infoset for that player. An important implication of this property is that each branch from the root to the leaves of the game tree traverses any given infoset at most once. We use the notation $z[I]$ to denote the history prefix of a terminal history $z$ that passes through infoset $I$. This notation is undefined (or invalid) if the path from $h^{0}$ to $z$ does not traverse $I$.
@@
$\pi_{i}^{\sigma}(h,h^{\prime})=\begin{cases}0,&\!\!\!\text{if }h\not\sqsubseteq h^{\prime}\\ \prod\limits_{h^{\prime\prime}:h\sqsubseteq h^{\prime\prime},h^{\prime\prime}\cdot a\sqsubseteq h^{\prime},\mathcal{P}(h^{\prime\prime})=i}\!\!\!\sigma_{i}(h^{\prime\prime},a),&\!\!\!\text{if }h\sqsubseteq h^{\prime}\end{cases}.$
@@
In game theory, the Nash equilibrium is a crucial solution concept: a state where no player can enhance their payoff by unilaterally altering their strategy, and to some extent the optimal solution. The payoff for player $i$ under strategy profile $\sigma$ is $u_{i}(\sigma)=\sum_{z\in Z}u_{i}(z)\pi^{\sigma}(z)$. The best response value for player $i$ against $\sigma_{-i}$ is $b_{i}(\sigma_{-i})=\max_{\sigma_{i}^{\prime}\in\Sigma_{i}}u_{i}(\sigma_{i}^{\prime},\sigma_{-i})$, the max payoff for optimal response. A strategy profile $\sigma^{*}$ is an $\epsilon$-Nash equilibrium if $\forall i\in\mathcal{N}\backslash\{c\}$, $u_{i}(\sigma^{*})+\epsilon\geq b_{i}(\sigma^{*}_{-i})$. When $\epsilon=0$, it’s a standard Nash equilibrium. In two-player zero-sum games, the exploitability of $\sigma$ is $\epsilon^{\sigma}=b_{1}(\sigma_{2})+b_{2}(\sigma_{1})$, measuring players’ vulnerability to an optimal opponent. Lower exploitability means closer to equilibrium.
@@ 2.1 Counterfactual Regret Minimization
Counterfactual Regret Minimization (CFR), also referred to as Vanilla CFR, is an iterative algorithm for computing approximate Nash equilibrium in IIEFGs by minimizing counterfactual regret (zinkevich2007regret). For player $i$, the total regret after $T$ iterations is defined as
@@
In two-player zero-sum IIEFGs, if both players satisfy $R^{T}_{i}\leq\epsilon$, then the average strategy forms a $2\epsilon$-Nash equilibrium (waugh2009abstraction).
@@
is used to update the immediate strategy via regret matching (hart2000simple):
$\sigma^{T+1}_{\mathcal{P}(I)}(I,a)=\begin{cases}\frac{R^{T}(I,a)_{+}}{~~\sum\limits_{\mathclap{a^{\prime}\in A(I)}}\left(R^{T}(I,a^{\prime})_{+}\right)},&\text{if }\sum\limits_{\mathclap{a^{\prime}\in A(I)}}\left(R^{T}(I,a^{\prime})_{+}\right)>0,\\ \frac{1}{|A(I)|},&\text{otherwise,}\end{cases}$
@@ 3.2 Driving Process for Embedding CFR
Building on this notation, we propose Embedding CFR, which models the extended information set abstraction problem as a strategy voting framework, where a embedding matrix $\mathbf{\Phi}=[\phi]_{m\times n}$ is constructed. Each non-negative element $\phi_{p,q}\geq 0$–with the constraint that $\sum_{p=1}^{m}\phi_{p,q}=1$ for each qq–represents the confidence in trusting advisor $e_{p}$ at infoset $I_{q}$; here, $\mathbf{\Phi}_{\cdot,q}$ denotes the embedding coordinates of infoset $I_{q}$. Given a strategy $\sigma_{i}^{T}(\mathbf{E})$ in the advisor space (also called embedding space), decisions in the original strategy space are made via the mapping
@@
To solve games under this abstraction, Embedding CFR operates as a no-regret learning algorithm that draws insights from Vanilla CFR (zinkevich2007regret): as illustrated in the right half of Figure 2 and as we have detailed in Section 2.1, Vanilla CFR updates the cumulative counterfactual regret $R^{T}(I)$ of the target infoset $I$ at iteration $T$, then derives the immediate strategy $\sigma_{i}^{T+1}(I)$ for next-iteration online interaction with the environment via regret matching, after which the regret for the new iteration is updated—forming a driving process. The strategy to be evaluated, which exhibits favorable convergence properties, is the average strategy $\bar{\sigma}_{i}^{T}(I)$. Embedding CFR extends this process by addressing how to drive it within the newly introduced advisor space.
@@
It is important to note that while regret embedding (Equation (5)) and strategy recovery (Equations (4), (8)) involve full matrix multiplications–specifically of sizes $m\times n$ with $n\times m$ and $n\times m$ with $m\times n$–this might initially suggest unavoidable $\mathcal{O}(n)$ or higher space complexity. Our solution addresses this by framing regret embedding as a sampling process and strategy recovery as a query procedure. Rather than operating on the full set of infosets, we dynamically engage a subset of $l<n$ infosets $I_{(1)},\dots,I_{(l)}$ (mapping to indices $j_{1},\dots,j_{l}$ in $\mathbf{J}$) at each iteration. As visualized in Figure 2, this selective focus is concretely illustrated: iteration $T$ activates only the single infoset $I_{4}$, while iteration $T+1$ shifts to $I_{3}$.
@@
An additional consideration is how to efficiently retrieve the embedding coordinates $\mathbf{\Phi}_{\cdot,j_{k}}$ for each infoset $I_{(k)}$–this will be discussed in the next section. Notably, each $\mathbf{\Phi}_{\cdot,j_{k}}$ only incurs $\mathcal{O}(m)$ space overhead, which does not significantly add to overall storage requirements. The blue and green arrows in Figure 2 illustrate these dynamic pathways, with purple nodes representing the quantities that need to be stored in this iteration—emphasizing that this paired sampling-and-query mechanism ensures Embedding CFR operates with $\mathcal{O}(m)$ space complexity. Pseudocode for poker-specific Embedding CFR is in Appendix B .